A company that creates education products is planning to create a suite of books to help customers prepare for entrance exams into college and graduate school. They have 33 in-house writers to create these books. Due to the expertise needed in creating this content it will not be possible to hire temporary writers within the planned time-frame. The projects are:# of Writers Needed NPV $800,000 $250,000 Test Prep Book Undergraduate Exam Prep Dental School Exam Prep Graduate School Exam

iven the resource constraint of only 33 available in-house writers, in which of these projects should your firm invest FIRST. Put another way, which of these projects is most attractive and efficient in the creation of value?

a.

Undergraduate Exam Prep

b.

Dental School Exam Prep

c.

Graduate School Exam Prep

d.

Law School Exam Prep

e.

Medical School Exam Prep
Answer #1

NPV/Number of Writers Needed

1.
Undergraduate=800000/24=33333.33

2.
Dental=250000/8=31250.00

3.
Graduate=450000/12=37500.00

4.
Law=320000/9=35555.56

5.
Medical=400000/11=36363.64

Graduate School Exam Prep is the most attractive and efficient in creation of value as it has the highest NPV/Number of Writers Needed